OpenStudy (anonymous):

The sum of two numbers is 51. Their difference is 13. What is the smaller of the two numbers? A. 18 B. 19 C. 20 D. 21

hero (hero):

x + y = 51 x - y = 13 Solve the system

OpenStudy (radar):

x+y=51 x-y=13 -------- 2x = 64 x=32 y=51-32=19

OpenStudy (radar):

Choice B